22 January 2008

RHIC Running:  The vernier scan did not run yesterday until 19:00.  Luminosities at STAR and PHENIX were higher after the scan than before.  The scan was performed on a store that had poor Yellow intensity because of an error in filling. 

Alex Zaltsman has requested an access to check on the L-10 cavity.  He believes that there may be a water leak in the piping, and is prepared to replace the pipe if a leak is found.   The repair will not take place during this morning's access.

Both experiments reported that even though beam has been erratic this week, they have recorded over 90% of the peak operating luminosity, so things are not that bad. 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020